Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News, Travel, Visitors (McMullin)

  • Overall status: B configuration; continued operational issues principally rooted in issues with the correlator back-end software.
  • Key dates:
    • 07 May - 15 Jun: 4-band science observing
  • Staff Travel:
  • Visitors:
    • 05 Jan - 06 Apr: John Cannon sabbatical at DSOC
    • 28 Feb - 28 Mar: Bob Sault

Hardware status and updates (Langley, Perley, Rupen)

  • Array (Perley/Langley):
  • WIDAR (Rupen):
    • Notes (Rupen/Sowinski): Good: Episodes with all four quadrants operating; 4x recirculation tests. Bad: Continued missing scans. Ugly: Memory issues (which may be linked to the missing scans) have led to significant instability, observing time loss.
  • Computing (Robnett):

Software status and updates (Butler)

  • Links:
  • Executor - updated CALC library to fix/improve planetary parallax calculations; update to MCAF to fix sideband issue identified last week. OPT tests on new production computer - chewbacca is ~3x faster.
  • CASA Update (new functionality available/issues/enhancements):Notes (Miriam): 3.3 planning complete and distributed.
  • AIPS Update (new functionality available/issues/enhancements): Notes (Eric): TYAPL fixed scaling (1.178); cleaned up integration times; TRUEP task is coming.
  • SDM Update: Notes:

Commissioning Group Reports and Test Planning

  • NOTE: Please cc me on submitted tests so that it can be assigned the appropriate priority!

Commissioning priorities for the week (McMullin, Chandler)

Single maintenance day week:
  • Mon
    • 1030-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Tue
    • %RED0800-1700: Testing/software development time
    • 1800-0630: Science operations and testing
  • Wed: Maintenance day
    • 0630-1700: Maintenance operations
    • 1800-0900: Science operations and testing
  • Thu, Fri:
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Sat-Sun
    • Science operations + planned testing

  • Plan to use Tuesdays/Friday as possible for science observing.

Issues Affecting Science Observing:

  • Correlator: Missing/mislabeled scans in data sets (small percentage but recurring). May be related to overall issues with CBE; major focus
  • Correlator: fshift sub-band anti-aliasing causing phantom spectral lines when using recirculation for narrow bands (only worry if Michael has to prep your file); Bob's report from last week; wobbles seem to be gone; Zeeman tests indicate another layer of problems.
  • Correlator: time mislabeling in SDM data (non-synchronous operation between correlator dumps and CBE can require additional 1/2 integration time padding for online flags);Update from last week; no further work due to other CBE issues; will resume once this is solved.
  • Array/Construction: Frequency offset tests (Dhawan/Rupen)
  • Correlator: Delay rocking across bands. Barry has written a proposal; pending implementation
  • Array: Amplitude wobbles seen on subset of antennas; phase jumps at scan boundaries also seen on subset of antennas (Ka, Q bands only thus far).

Commissioning Groups

Testing This week

  • Correlator: Rupen: Subarray testing; CBE merge??
  • Solar: Perley/Bastian: Solar observing tests (20dB testing); Software control of attenuator now available.
  • Spectral line: Momjian: HI High resolution bandpass testing
  • Spectral line: Momjian: Zeeman test follow-up

RSRO/ECSO Reports

RSRO B configuration status: https://safe.nrao.edu/wiki/bin/view/EVLA/2011BRSROStatus

Individual reports (updates)

Misc:

Demo Science Status

AOB

  • Looking at a public web-based observing log (searchable) (Butler/McMullin).

Misc Actions

  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • Development of WIDAR guidelines document (Michael)
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).

-- JosephMcMullin - 2011-03-21
Topic revision: r3 - 2011-03-29, JosephMcMullin
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback